What is the molality of para-dichlorobenzene in a solution prepared by dissolving 2.65g C6H4Cl2 in 50.0 mL of benzene (d=0.897g/mL)?

1
Expert's answer
2022-05-30T07:52:02-0400

Solution:

The solute is para-dichlorobenzene (C6H4Cl2)

The solvent is benzene (C6H6)


Density = Mass / Volume

Therefore,

Mass of C6H6 = Density × Volume = (0.897 g mL−1) × (50.0 mL) = 44.85 g

Kilograms of C6H6 = (44.85 g) × (1 kg / 1000 g) = 0.04485 kg


Moles = Mass / Molar mass

The molar mass of para-dichlorobenzene (C6H4Cl2) is 147 g mol−1

Therefore,

Moles of C6H4Cl2 = (2.65 g) / (147 g mol−1) = 0.01803 mol


Molality = Moles of solute / Kilograms of solvent

Therefore,

Molality of C6H4Cl2 solution = Moles of C6H4Cl2 / Kilograms of C6H6

Molality of C6H4Cl2 solution = (0.01803 mol) / (0.04485 kg) = 0.402 mol kg−1 = 0.402 m

Molality of C6H4Cl2 solution = 0.402 m


Answer: The molality of para-dichlorobenzene (C6H4Cl2) solution is 0.402 m